# TextCenGen: Attention-Guided Text-Centric Background Adaptation for Text-to-Image Generation[Uncaptioned image]

### v4

@@ Abstract
Recent advancements in Text-to-Image (T2I) generation have witnessed a shift from adapting text to fixed backgrounds to creating images around text. Traditional approaches are often limited to generate layouts within static images for effective text placement. Our proposed approach, TextCenGen, introduces a training-free and plug-and-play dynamic background adaptation of the blank region for text-friendly image generation. Our method employs force-directed attention guidance in T2I models to generate images that strategically reserve whitespace for any pre-defined text areas. Observing how cross-attention maps affect object placement, we detect and repel conflicting objects using a force-directed graph approach, combined with a spatial excluding cross-attention constraint for smooth attention in whitespace areas. To evaluate our method, we propose hybrid prompt datasets as benchmarks with . Experiments indicate that TextCenGen outperforms existing methods by producing more harmonious compositions, demonstrating that it effectively enhances T2I model outcomes by catering to different text positions.
@@
Figure 1: TextCenGen is a training-free method designed to generate text-friendly images. By using a simple text prompt and a planned blank region as inputs, TextCenGen creates images that satisfy the prompt and provide sufficient blank space in the target region. For example, the text-friendly T2I approach helps users customize their favored text-friendly wallpapers for mobile devices with any T2I model, avoiding visual confusion caused by main objects overlapping with UI components.
Introduction
111*These authors contributed equally to this work222†Corresponding author: chli@cs.ecnu.edu.cn
@@ Introduction
In graphic design, achieving an effective synergy between text and imagery is essential for clear communication. The choice of background can influence text visibility and comprehension. A common design objective is to place text within an image in a way that is both visually pleasing and communicatively clear. A preferred strategy is to position the text in the golden ratio, which is believed to be aesthetically optimal. However, designers often grapple with the issue of backgrounds that compete with or obscure the text, as illustrated in Figure 1[Uncaptioned image], where unsuitable backgrounds detract from the text’s readability and aesthetic appeal, regardless of any adjustments to text color or size. We aim to facilitate the creation of text-friendly images (see Figure 1[Uncaptioned image]), ideal for text placement and meeting the growing demand due to increasing use of Text-to-Image (T2I) models for background graphics.
@@
Traditional approaches to graphic design, especially in poster creation, have largely focused on arranging layouts in static natural background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). However, producing text-friendly images remains a significant challenge due to the complexity of background elements. Our insight, derived from Figure 1, reveals that a clear separation between the main objects and the text areas is essential. Recent advancements in diffusion-related research have shown that it is possible to manipulate the primary objects within cross-attention maps, making the adaptation of background images to accommodate text a feasible endeavor (Hertz et al. 2022; Epstein et al. 2023).
@@
In response, we introduce TextCenGen[Uncaptioned image], as illustrated in Figure 1, a novel method that employs cross-attention maps and force-directed graphs for effective object placement and whitespace optimization. We also implement a spatial excluding cross-attention constraint to ensure smooth attention in areas designated for text. To establish a new benchmark for this innovative task, we constructed a diverse dataset gathered from three unique sources, along with three distinct evaluation metrics, to comprehensively assess the performance. The contributions of our paper are three-fold:
@@
  • We propose a new task of text-freindly T2I generation, which creates images that satisfy both the prompt and reserve space for pre-defined text placements. The task consists of a benchmark including a specialized dataset and tailored evaluation metrics.
@@
  • We introduce TextCenGen, a plug-and-play, training-free framework. This approach allows for dynamic placement of text at user-specified location within an image, with the image being generated or adapted to accommodate the text.
  • Our prominent module, the force-directed cross-attention guidance, is an effective addition to the denoising process in T2I models. This module strategically directs cross-attention map within the image generation process, ensuring harmonious layout of text and imagery.
Related Work
@@
Text Layout of Natural Images has evolved significantly, transitioning from traditional layout designs to more advanced methods influenced by deep learning. Initially, poster design focused on creating layouts with given background images, elements, and text (Guo et al. 2021; Cao, Chan, and Lau 2012; O’Donovan, Agarwala, and Hertzmann 2014; Li, Zhang, and Wang 2022). The integration of deep learning in text layout of natural image generation has led to various models such as GAN (Goodfellow et al. 2014; Zheng et al. 2019; Li et al. 2019; Zhou et al. 2022), VAE (Jyothi et al. 2019), transformers (Vaswani et al. 2017; Inoue et al. 2023b; Wang et al. 2023) and diffusion models (Ho, Jain, and Abbeel 2020; Hui et al. 2023; Inoue et al. 2023a; Chai, Zhuang, and Yan 2023; Li et al. 2023a). These models have been instrumental in learning layout patterns from large datasets. Subsequent research explored methods to retrieve matching background images based on text and image description (Jin et al. 2022). Since the development of image editing method based on diffusion model, scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and DiffText (Zhang et al. 2023) have addressed the challenges of generating clear text with diffusion models. However, these methods often rely on the presence of a “sign” or similar element within the prompt (e.g., a T-shirt) to place text. They do not explicitly tackle the problem of creating text-friendly images where the background itself is crafted to adapt pre-defined text regions. Our approach extends these capabilities by allowing the primary objects in generated images to yield space to text regions, resulting in more harmonious and aesthetically pleasing compositions.
@@
Text-to-Image Generation has advanced with diffusion models (Ho, Jain, and Abbeel 2020), producing realistic images and videos that align closely with text prompts (Rombach et al. 2022; Singer et al. 2023; Chen et al. 2023c). Innovations in this field include GLIDE (Nichol et al. 2022), which integrates text conditions into the diffusion process, Dall-E 2 (Ramesh et al. 2022) with its diffusion prior module for high-resolution images, and Imagen (Saharia et al. 2022), which uses a large T5 language model to enhance semantic representation. Stable diffusion (Rombach et al. 2022) projects images into latent space for diffusion processing. Beyond text conditions, the manipulation of diffusion models through image-level conditions has been explored. Methods such as image inpainting (Balaji et al. 2022) aim to generate coherent parts of an image, while SDG (Liu et al. 2023) introduces semantic inputs to guide unconditional DDPM sampling. In addition, techniques such as (Meng et al. 2021) use images as editing conditions in denoising processes. Scene text generation methods such as TextDiffuser (Chen et al. 2023b, a) and GlyphDraw (Ma et al. 2023) have also emerged, utilizing textual layouts or masks to guide text generation in images. These advancements represent the growing versatility and potential of T2I models in diverse applications.
@@
Attention Guided Image Editing has emerged as a foundamental solution to the challenge of translating human preferences and intentions into visual content through text descriptions. These approaches, as seen in works like (Li et al. 2023b; Avrahami et al. 2023; Zhang, Rao, and Agrawala 2023; Zhao et al. 2023), involve learning auxiliary modules on paired data. However, a significant limitation of these training-based methods is the substantial cost and effort required for repeated training for different control signals, model architectures, and checkpoints. In response to these challenges, training-free techniques have emerged, using the inherent weights of attention and the characteristics of pre-trained models to control the characteristics of objects such as size, shape, appearance, and location (Hertz et al. 2022; Epstein et al. 2023; Patashnik et al. 2023; Xie et al. 2023). These methods typically utilize basic conditions, such as bounding boxes, for precise control over object positioning and scene composition (Mo et al. 2023). Desigen (Weng et al. 2024) discovered relationship between attention and saliency and introduce attention reduction to weaken the attention whin layout boxes. Our approach takes this further by applying force-directed graph techniques to cross-attention map edits, allowing for more automated and precise object transformations in T2I editing.
Method
@@
Given an input text region $R$ and a prompt $T_{prompt}$, our framework aims to generate a text-friendly image $I_{res}$. Specifically, the output image has (1) reduced overlap between the primary object and $R$, (2) sufficient smoothness and minimal color variation in $R$, and (3) the image still fits the prompt. Our framework is shown in Figure 2.
As cross-attention map $A_{k}$ serves as a medium to locate and edit objects within generated images (Epstein et al. 2023), we focus on the $k^{th}$ token of $T_{prompt}$. In response to our first concern, we analyze the denoising process of an unguided image ($I_{ori}$) from a given prompt, and establish the subset of tokens ($O$) that refer to objects that need modification. We design a conflict detector that determines object conflicts based on the average attention intensity in the overlapping regions between $A_{k}$ and $R$. For every token $k\in O$, we introduce Force-Directed Cross-Attention Guidance for moving objects. In this scheme, objects are treated as centroid vertex within a graph, with a sequence of forces being applied (see Figure 3) to adjust object positions.
For the second consideration, inspired by recent technologies that limit the range of the attention map (Zhang et al. 2023), we propose a spatial excluding cross-attention constraint to prevent an extensive attention density from encroaching on $R$.
@@
Force-Directed Cross-Attention Guidance
Figure 3: Illustration of four set relationships and their associated forces. The Repulsive Force separates object and text centroids during intersections (a1) and enclosures (a2). The Margin Force (b) and Warping Force (c) prevent boundary overstepping. Text within object regions (a4) requires cooperation between force and attention constraint.
@@
To seamlessly integrate the concept of force-directed graphs into the loss guidance of the denoising process in latent diffusion models, we delve into the extraction and manipulation of attention maps and activations. We use softmax normalized attention matrices $\mathcal{A}_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times K}$ of layer $i$ extracted from the standard denoising forward step $\epsilon_{\theta_{i}}(z_{i};t,y)$. This enables us to manipulate the control over objects referred to in the text conditioning $y$ at distinct indices $k$, by adjusting the related attention channel(s) $\mathcal{A}_{i,t,\ldots,k}\in\mathbb{R}^{H_{i}\times W_{i}\times|k|}$. The centroid of the attention map is a two-dimensional vector, defined by the equation:
@@
To ensure the efficacy of our centroid-based object representation method, we initially assume that all objects are convex sets, adhering to the mathematical definition that for every pair of points within the object, the line segment connecting them lies entirely within the object. This assumption allows us to treat the extracted $\text{centroid}\left(k\right)$ as vertices $v_{k}$ in a graph, which are then subjected to force-directed attention guidance. Indeed, it is important to clarify that while each token $k$ associated with $\mathcal{A}_{k}$ appears as a single entity in Figure 2, it actually represents an average of $\mathcal{A}_{k}^{l}$ across all layers $l$ in the U-Net architecture. Practically, our method is applied individually to each layer, ensuring a nuanced and layer-specific approach to force-directed attention guidance.
@@
$F_{multi-tar}(v_{i})=\sum_{j=1}^{n}\omega_{j}\cdot\frac{-\xi^{2}}{||p(v_{i})-p(tar_{j})||},$ (3)
@@
This matrix transforms the coordinates of $O$, ensuring that it remains visible within $A$ after transformation. This carefully planned process secures the region $O$ within $A$, even after dynamic changes. It upholds the structure of the cross-attention map and the core principles of our force-directed graph algorithm, balancing key vertices visibility and graph fluidity.
Spatial Excluding Cross-Attention Constraint

### v5

# Tianyi Liang    Jiangqi Liu    Yifei Huang    Shiqi Jiang    Jianshen Shi    Changbo Wang    Chenhui Li

@@ Abstract
Text-to-image (T2I) generation has made remarkable progress in producing high-quality images, but a fundamental challenge remains: creating backgrounds that naturally accommodate text placement without compromising image quality. This capability is non-trivial for real-world applications like graphic design, where clear visual hierarchy between content and text is essential. Prior work has primarily focused on arranging layouts within existing static images, leaving unexplored the potential of T2I models for generating text-friendly backgrounds. We present TextCenGen, a training-free dynamic background adaptation in the blank region for text-friendly image generation. Instead of directly reducing attention in text areas, which degrades image quality, we relocate conflicting objects before background optimization. Our method analyzes cross-attention maps to identify conflicting objects overlapping with text regions and uses a force-directed graph approach to guide their relocation, followed by attention excluding constraints to ensure smooth backgrounds. Our method is plug-and-play, requiring no additional training while well balancing both semantic fidelity and visual quality. Evaluated on our proposed text-friendly T2I benchmark of 27,000 images across four seed datasets, TextCenGen outperforms existing methods by achieving 23% lower saliency overlap in text regions while maintaining 98% of the semantic fidelity measured by CLIP score and our proposed Visual-Textual Concordance Metric (VTCM).
Machine Learning, ICML
@@
Figure 1: TextCenGen is a training-free method designed to generate text-friendly images. By using a simple text prompt and a planned blank region as inputs, TextCenGen creates images that satisfy the prompt and provide sufficient blank space in the target region. For example, the text-friendly T2I approach helps users customize their favored text-friendly wallpapers for mobile devices with any T2I model, avoiding visual confusion caused by the main objects overlapping with UI components.
1 Introduction
@@
In graphic design, achieving a harmonious visual effect between text and imagery is essential for clear expression. The choice of background can influence text visibility and comprehension. A common design objective is to place text within an image in a way that is both visually pleasing and clearly conveys the intended message. A preferred strategy is to position the text in the golden ratio, which is believed to be aesthetically optimal. However, designers often grapple with the issue of backgrounds that compete with or obscure the text, as illustrated in Figure 1[Uncaptioned image], where unsuitable backgrounds detract from the text’s readability and aesthetic appeal, regardless of any adjustments to text color or size. We aim to facilitate the creation of text-friendly images (see Figure 1[Uncaptioned image]), ideal for text placement and meeting the growing demand due to the increasing use of Text-to-Image (T2I) models for background graphics.
@@
Traditional approaches to graphic design, especially in poster creation, have largely focused on arranging layouts in static natural background images, elements, and text (Guo et al., 2021; Cao et al., 2012; O’Donovan et al., 2014; Li et al., 2022). However, producing text-friendly images remains a challenge due to the complexity of background elements. Our insight, derived from Figure 1, reveals that a clear separation between the main objects and the text areas is essential. Recent advancements in diffusion-related research have shown that it is possible to manipulate the primary objects within cross-attention maps, making the adaptation of background images to accommodate text a feasible endeavor (Hertz et al., 2022; Epstein et al., 2023; Wang et al., 2024). However, we discovered that directly reducing attention in the target region leads to a semantic reduction in the generated image’s match with the prompt. Could we move the conflicting objects out of the target area before reducing attention?
@@
In response, we introduce TextCenGen[Uncaptioned image] [^1] Open source code at: https://github.com/tianyilt/TextCenGen_Background_Adapt, as illustrated in Figure 1, a new method that employs cross-attention maps and force-directed graphs for effective object placement and whitespace optimization. We also implement a spatial excluding cross-attention constraint to ensure smooth attention in areas designated for text. To establish a new benchmark for this innovative task, we constructed a diverse dataset gathered from three unique sources, along with five evaluation metrics, to comprehensively assess the performance. The contributions of our paper are three-fold:
@@
  • We propose a new task of text-friendly T2I generation, which creates images that satisfy both the prompt and reserve space for pre-defined text placements. The task consists of a benchmark including a specialized dataset and tailored evaluation metrics.
@@
  • We introduce TextCenGen, a plug-and-play, training-free background adaptation framework for dynamic text placement in generated images.
  • We develop force-directed cross-attention guidance, adaptable to various attention mechanisms across different T2I models, ensuring a harmonious layout of text and imagery.
2 Related Work
@@
Text Layout of Natural Images has evolved significantly, transitioning from traditional layout designs to more advanced methods influenced by deep learning. Initially, poster design focused on creating layouts with given background images, elements, and text (Guo et al., 2021; Cao et al., 2012; O’Donovan et al., 2014; Li et al., 2022; Zhang et al., 2020). The integration of deep learning in text layout of natural image generation has led to various models such as GAN (Goodfellow et al., 2014; Zheng et al., 2019; Li et al., 2019; Zhou et al., 2022), VAE (Jyothi et al., 2019), transformers (Vaswani et al., 2017; Inoue et al., 2023a; Wang et al., 2023) and diffusion models (Ho et al., 2020; Hui et al., 2023; Inoue et al., 2023b; Chai et al., 2023; Li et al., 2023a). These models have been instrumental in learning layout patterns from large datasets. Subsequent research explored methods to retrieve matching background images based on text and image description (Jin et al., 2022). Since the development of image editing method based on diffusion model, scene text generation methods such as TextDiffuser (Chen et al., 2023a, b) and DiffText (Zhang et al., 2024) have addressed the challenges of generating clear text with diffusion models. However, these methods often rely on the presence of a “sign” or similar element within the prompt (e.g., a T-shirt) to place text. They do not explicitly tackle the problem of creating text-friendly images where the background itself is crafted to adapt pre-defined text regions. Our approach extends these capabilities by allowing the primary objects in generated images to yield space to text regions, resulting in more harmonious and aesthetically pleasing compositions.
@@
Text-to-Image Generation has advanced with diffusion models (Ho et al., 2020), producing realistic images and videos that align closely with text prompts (Rombach et al., 2022; Singer et al., 2023; Chen et al., 2023c; Esser et al., 2024). Innovations in this field include GLIDE (Nichol et al., 2022), which integrates text conditions into the diffusion process, Dall-E 2 (Ramesh et al., 2022) with its diffusion prior module for high-resolution images, and Imagen (Saharia et al., 2022), which uses a large T5 language model to enhance semantic representation. Stable diffusion (Rombach et al., 2022) projects images into latent space for diffusion processing. Beyond text conditions, the manipulation of diffusion models through image-level conditions has been explored. Methods such as image inpainting (Balaji et al., 2022) aim to generate coherent parts of an image, while SDG (Liu et al., 2023) introduces semantic inputs to guide unconditional DDPM sampling. In addition, techniques such as (Meng et al., 2021) use images as editing conditions in denoising processes. Scene text generation methods such as TextDiffuser (Chen et al., 2023a, b) and GlyphDraw (Ma et al., 2023) have also emerged, utilizing textual layouts or masks to guide text generation in images. These advancements represent the growing versatility and potential of T2I models in diverse applications.
@@
Attention Guided Image Editing has emerged as a fundamental solution to the challenge of translating human preferences and intentions into visual content through text descriptions. These approaches, as seen in works like (Li et al., 2023b; Avrahami et al., 2023; Zhang et al., 2023; Zhao et al., 2023), involve learning auxiliary modules on paired data. However, a limitation of these training-based methods is the substantial cost and effort required for repeated training for different control signals, model architectures, and checkpoints. In response to these challenges, training-free techniques have emerged, using the inherent weights of attention and the pre-trained models to control the attributes of objects such as size, shape, appearance, and location (Hertz et al., 2022; Epstein et al., 2023; Patashnik et al., 2023; Xie et al., 2023; Zhou et al., 2024a, b). These methods typically utilize basic conditions, such as bounding boxes, for precise control over object positioning and scene composition (Mo et al., 2023). Desigen (Weng et al., 2024) discovers relationship between attention and saliency and introduces attention reduction to weaken the attention within layout boxes. Our approach takes this further by applying force-directed graph techniques to cross-attention map edits, allowing for more automated and precise object transformations in T2I editing.
3 Method
@@
Given an input text region $R$ and a prompt $T_{prompt}$, our framework aims to generate a text-friendly image $I_{res}$. This research is motivated by extensive interviews with wallpaper designers, who emphasized the need for precise control over text areas, object positioning, and background consistency in T2I-model-generated images. Specifically, our framework addresses these concerns by producing an output image that has (1) reduced overlap between the primary object and $R$, (2) sufficient smoothness and minimal color variation in $R$, and (3) the image still fits the prompt. Our framework is shown in Figure 2.
As cross-attention map $A_{k}$ serves as a medium to locate and edit objects within generated images (Epstein et al., 2023), we focus on the $k^{th}$ token of $T_{prompt}$. In response to our first concern, we analyze the denoising process of an unguided image ($I_{ori}$) from a given prompt, and establish the subset of tokens ($O$) that refer to objects that need modification. We design a conflict detector that determines object conflicts based on the average attention intensity in the overlapping regions between $A_{k}$ and $R$. For every token $k\in O$, we introduce Force-Directed Cross-Attention Guidance for moving objects. In this scheme, objects are treated as centroid vertex within a graph, with a sequence of forces being applied (see Figure 3) to adjust object positions.
For the second consideration, inspired by recent technologies that limit the range of the attention map (Zhang et al., 2024), we propose a spatial excluding cross-attention constraint to prevent an extensive attention density from encroaching on $R$.
@@
3.1 Force-Directed Cross-Attention Guidance
Figure 3: Illustration of four set relationships and their associated forces. The Repulsive Force separates object and text centroids during intersections (a1) and object in text (a2). The Margin Force (b) and Warping Force (c) prevent boundary overstepping. Text within object regions (a4) requires cooperation between force and attention constraint. Separation (a3) isn’t required to process.
@@ 3.1 Force-Directed Cross-Attention Guidance
To seamlessly integrate the concept of force-directed graphs into the loss guidance of the denoising process in latent diffusion models, we delve into the extraction and manipulation of attention maps and activations. For denoising image $i$, we use softmax normalized attention matrices $\mathcal{A}_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times K}$ extracted from the standard denoising forward step $\epsilon_{\theta_{i}}(z_{i};t,y)$. This enables us to manipulate the control over objects referred to in the text conditioning $y$ at distinct indices $k$, by adjusting the related attention channel(s) $\mathcal{A}_{i,t,\ldots,k}\in\mathbb{R}^{H_{i}\times W_{i}\times|k|}$. The centroid of the attention map is a two-dimensional vector, defined by the equation:
@@
We assume that all objects are convex sets, adhering to the mathematical definition that for every pair of points within the object, the line segment connecting them lies entirely within the object. This assumption allows us to treat the extracted $\text{centroid}\left(k\right)$ as vertices $v_{k}$ in a graph, which are then subjected to force-directed attention guidance. Indeed, it is important to clarify that while each token $k$ associated with $\mathcal{A}_{k}$ appears as a single entity in Figure 2, it actually represents an average of $\mathcal{A}_{k}^{l}$ across all layers $l$ in the U-Net architecture. Practically, our method is applied individually to each layer, ensuring a nuanced and layer-specific approach to force-directed attention guidance.
@@
$F_{mt}(v_{i})=\sum_{j=1}^{n}\omega_{j}\cdot\frac{-\xi^{2}}{||p(v_{i})-p(tar_{j})||},$ (3)
@@
This matrix transforms the coordinates of $O$, ensuring that it remains visible within $A$ after transformation. This carefully planned process secures the region $O$ within $A$, even after dynamic changes. It upholds the structure of the cross-attention map, balancing key vertices visibility and graph fluidity.
Dataset Metrics Dall-E 3 AnyText Desigen SD 1.5 Ours (SD 1.5) Improve (%)
P2P Template Saliency IOU $\downarrow$ 52.64 54.34 30.62 29.89 $\mathbf{22.86}$ 23.52% $\uparrow$
TV Loss $\downarrow$ 18.02 22.55 13.7 14.11 $\mathbf{8.81}$ 37.56% $\uparrow$
VTCM $\uparrow$ 1.92 1.75 2.92 2.95 $\mathbf{4.4}$ 49.15% $\uparrow$
DiffuisonDB Saliency IOU $\downarrow$ 56.82 55.88 30.3 30.11 $\mathbf{23.59}$ 21.65% $\uparrow$
TV Loss $\downarrow$ 21.16 20.39 11.99 12.3 $\mathbf{8.41}$ 31.63% $\uparrow$
VTCM $\uparrow$ 1.67 1.79 3.20 3.19 $\mathbf{4.39}$ 37.62% $\uparrow$
Syn Prompt Saliency IOU $\downarrow$ 51.52 53.24 31.57 31.42 $\mathbf{27.7}$ 11.84% $\uparrow$
TV Loss $\downarrow$ 17.85 21.46 15.63 15.61 $\mathbf{11.37}$ 27.16% $\uparrow$
VTCM $\uparrow$ 1.98 1.79 2.67 2.73 $\mathbf{3.49}$ 27.84% $\uparrow$
Table 1: Quantitative comparison of metrics across different methods and datasets. Bold indicate the best scores.
Dataset Metrics SD 1.5 Ours 1.5 Imp (%) SD 2.0 Ours 2.0 Imp (%) SDXL Ours XL Imp (%)
P2P Saliency IOU $\downarrow$ 29.89 $\mathbf{22.86}$ 23.52 37.97 33.33 12.22 29.83 26.64 10.69
Template TV Loss $\downarrow$ 14.11 $\mathbf{8.81}$ 37.56 17.73 12.06 31.98 12.09 9.10 24.73
VTCM $\uparrow$ 2.95 $\mathbf{4.40}$ 49.15 2.52 3.38 34.13 3.41 4.24 24.34
DiffusionDB Saliency IOU $\downarrow$ 30.11 $\mathbf{23.59}$ 21.65 33.40 29.52 11.62 34.22 32.31 3.16
TV Loss $\downarrow$ 12.30 $\mathbf{8.41}$ 31.63 16.17 12.78 20.96 13.22 10.66 15.86
VTCM $\uparrow$ 3.19 $\mathbf{4.39}$ 37.62 2.63 4.06 54.37 3.27 3.43 4.89
Syn Saliency IOU $\downarrow$ 31.42 $\mathbf{27.70}$ 11.84 38.59 36.22 6.14 28.84 24.77 14.11
Prompt TV Loss $\downarrow$ 15.61 $\mathbf{11.37}$ 27.16 18.92 15.25 19.40 12.31 8.48 31.11
VTCM $\uparrow$ 2.73 $\mathbf{3.49}$ 27.84 2.42 2.83 16.94 3.59 4.78 33.15
Table 2: Performance comparison across different Stable Diffusion versions. Bold indicates the best scores for SD 1.5, which serves as our primary baseline. "Imp" represents the percentage improvement over the corresponding base model.
 Dataset Dall-E 3 AnyText Desigen Ours
P2P Template 2.53 0.32 0.55 $\mathbf{0.30}$
DiffuisonDB 2.04 1.04 0.34 $\mathbf{0.31}$
Syn Prompt 2.25 1.02 0.51 $\mathbf{0.29}$
Dataset Metrics
Desigen-TF +AR
Desigen-T +AR
Ours (SD 1.5)
P2P Template Saliency IOU $\downarrow$ 30.62 34.99 $\mathbf{22.86}$
TV Loss $\downarrow$ 13.70 14.77 $\mathbf{8.81}$
VTCM $\uparrow$ 2.92 2.97 $\mathbf{4.40}$
DiffusionDB Saliency IOU $\downarrow$ 30.30 31.11 $\mathbf{23.59}$
TV Loss $\downarrow$ 11.99 14.58 $\mathbf{8.41}$
VTCM $\uparrow$ 3.20 2.82 $\mathbf{4.39}$
Syn Prompt Saliency IOU $\downarrow$ 31.57 32.60 $\mathbf{27.70}$
TV Loss $\downarrow$ 15.63 14.51 $\mathbf{11.37}$
VTCM $\uparrow$ 2.67 2.93 $\mathbf{3.49}$
Table 4: Comparison with trained versions of Desigen on general datasets. Desigen-TF: Training-free, Desigen-T: Trained, AR: Attention Reduction. Even with specialized training on graphic design data, Desigen-T does not match our training-free method’s performance.
Method Saliency IOU $\downarrow$ TV Loss $\downarrow$ VTCM $\uparrow$
D-TF 40.79 19.44 2.36
D-T 41.66 18.06 2.30
D-T+Ours 38.48 12.19 2.85
Ours (SD 1.5) $\mathbf{31.99}$ $\mathbf{9.74}$ $\mathbf{3.69}$
Table 5: Performance on the specialized Desigen benchmark dataset. D-TF: Desigen-Training-free, D-T: Desigen-Trained. Underline indicates second-best performance.
Dataset Metrics Text2Poster Best@5 Text2Poster Avg@5 Ours (SD 1.5)
P2P Template Saliency IOU $\downarrow$ 31.25 36.22 $\mathbf{22.86}$
TV Loss $\downarrow$ 11.49 12.45 $\mathbf{8.81}$
VTCM $\uparrow$ 2.48 2.28 $\mathbf{4.40}$
Clip Score $\uparrow$ 20.80 20.80 $\mathbf{27.96}$
DiffusionDB Saliency IOU $\downarrow$ 24.07 34.38 $\mathbf{23.59}$
TV Loss $\downarrow$ $\mathbf{7.99}$ 10.65 8.41
VTCM $\uparrow$ 2.93 2.22 $\mathbf{4.39}$
Clip Score $\uparrow$ 17.57 17.57 $\mathbf{27.20}$
Syn Prompt Saliency IOU $\downarrow$ 31.25 36.91 $\mathbf{27.70}$
TV Loss $\downarrow$ 11.49 13.28 $\mathbf{11.37}$
VTCM $\uparrow$ 2.48 2.16 $\mathbf{3.49}$
Clip Score $\uparrow$ 20.80 20.90 $\mathbf{28.10}$
Table 6: Comparison with retrieval-based methods. Text2Poster Best@5 shows the best results selected from five preset positions, while Avg@5 shows the average. Our method outperforms Text2Poster in most metrics, particularly CLIP Score, demonstrating better semantic fidelity while maintaining text-friendliness.
3.2 Spatial Excluding Cross-Attention Constraint